# PRISM: A Robust Framework for Skill-based Meta-Reinforcement Learning with Noisy Demonstrations

### v2

@@ 4.1 Motivation: Challenges in Direct Skill Learning with Noisy Demonstrations
Figure 2: Comparison of skill learning in the microwave-opening task using expert and noisy demonstrations: (a) Learned skills, (b) Meta-RL performance utilizing the learned skills
Existing skill-based meta-RL methods, as discussed in Section 3, rely on offline datasets assumed to contain expert demonstrations. However, real-world datasets often suffer from action noise due to aging robots, environmental perturbations, or sensor errors. Unlike online setups that adapt through re-training, static offline datasets are vulnerable to noise. This significantly hinders skill learning, particularly in long-horizon tasks with cumulative errors and manipulation tasks requiring precision. Fig. 2(a) compares skills learned from expert and noisy datasets for a microwave-opening task in the Kitchen environment. Skills from the expert dataset successfully complete the task, while those from noisy demonstrations fail to grasp or open the microwave. Fig. 2(b) shows that SiMPL, a skill-based meta-RL method, performs poorly on test tasks when relying on skills learned from noisy data compared to expert data. This stems from existing methods learning indiscriminately from offline datasets, regardless of trajectory quality. To tackle the challenges posed by noisy offline datasets, we propose prioritized skill refinement, integrating online exploration and maximum return relabeling to dynamically identify and refine task-relevant skills. The following sections detail how the proposed methods leverage these components to ensure robust skill learning for meta-RL under noisy conditions.
Figure 3: Structure of the proposed PRISM framework
4.2 Prioritized Skill Refinement with Exploration
In this section, we introduce the prioritized refinement for skill-based meta-RL (PRISM) framework to address the limitations of traditional skill learning, which relies heavily on offline datasets often containing noisy and irrelevant trajectories, as shown in Fig. 2(a). To tackle this inefficiency, PRISM incorporates an exploration policy $\pi_{\mathrm{exp}}(a_{t}|s_{t},i)$ for each training task $\mathcal{T}_{i}$, where $i=1,\cdots,N_{\mathcal{T},\mathrm{train}}$, and $N_{\mathcal{T},\mathrm{train}}$ denotes the number of training tasks in $\mathcal{M}_{\mathrm{train}}$. While research on improving exploration capabilities has been conducted across various domains (Ermolov & Sebe, 2020; Han & Sung, 2021; Chu et al., 2024; Jo et al., 2024), we utilize intrinsic motivation techniques (Burda et al., 2018) to identify task-relevant trajectories by performing exploration near the noisy offline dataset initially. However, the abundance of irrelevant trajectories in the noisy offline dataset can hinder efficient learning.
To address this, PRISM introduces two key buffers: the exploration buffer $\mathcal{B}_{\text{exp}}^{i}$ and the prioritized online buffer $\mathcal{B}_{\text{on}}^{i}$. The exploration buffer $\mathcal{B}_{\text{exp}}^{i}$ stores all trajectories collected by the exploration policy, providing comprehensive samples necessary for RL training, defined as $\mathcal{B}_{\text{exp}}^{i}:=\{\tau_{\text{exp}}^{i}\}$, where $\tau_{\text{exp}}^{i}=(s_{0}^{i},a_{0}^{i},r_{0}^{i},\cdots,s_{H}^{i})$ represents an exploration trajectory, and $a_{t}^{i}\sim\pi_{\text{exp}}(\cdot|s_{t},i)$ for $t=0,\cdots,H$. The prioritized online buffer $\mathcal{B}_{\text{on}}^{i}$, initialized with the offline dataset $\mathcal{B}_{\text{off}}$, selectively retains high-quality trajectories based on their returns, defined as $\mathcal{B}_{\text{on}}^{i}:=\{\tau_{\text{high}}^{i}\}$, where $\tau_{\text{high}}^{i}$ are trajectories prioritized by their highest returns $G=\sum_{t=0}^{H-1}\gamma^{t}r_{t}^{i}$. These high-return trajectories include both exploration trajectories $\tau_{\text{exp}}^{i}$ collected by the exploration policy and low-level trajectories $\tau_{l}^{i}$ generated by the low-level policy $\pi_{l}(a_{t}|s_{t},z)$ during the execution of the high-level policy $\pi_{h}$, where $z\sim\pi_{h}(\cdot|s_{t},e^{\mathcal{T}_{i}})$ and $e^{\mathcal{T}_{i}}\sim q_{e}$.
By filtering and retaining only the most useful trajectories, $\mathcal{B}_{\text{on}}^{i}$ effectively guides the exploration policy, surpassing the noisy offline buffer in efficiency and utility, as inspired by self-supervised learning approaches (Xin et al., 2020; Kim et al., 2024). The exploration policy $\pi_{\text{exp}}$ is subsequently updated to minimize the following loss function, enabling it to focus on high-quality samples stored in $\mathcal{B}_{\text{on}}^{i}$ and discover additional task-relevant action sequences:
$\sum_{i}\mathbb{E}_{\tau^{i}\sim\mathcal{B}_{\text{exp}}^{i}\cup\mathcal{B}_{\text{on}}^{i}}\left[\mathcal{L}^{\text{RL}}_{\text{exp}}(\pi_{\text{exp}})\right]+\lambda^{\text{kld}}_{\text{exp}}\mathbb{E}_{\tau^{i}\sim\mathcal{B}_{\text{on}}^{i}}\mathcal{D}_{\text{KL}}(\hat{\pi}_{d}^{i}||\pi_{\text{exp}}),\vspace{-.5em}$ (3)
where $\lambda^{\text{kld}}_{\text{exp}}$ is the KLD coefficient for $\pi_{\text{exp}}$, and $\hat{\pi}_{d}^{i}$ denotes the action data distribution derived from $\mathcal{B}_{\text{on}}^{i}$. This setup ensures that the exploration policy focuses on task-relevant trajectories, starting near the offline dataset since the prioritized online buffer is initialized with it. Over time, as high-quality trajectories accumulate in the online buffer, the exploration policy progressively improves its ability to discover better paths and task-relevant actions.
Furthermore, PRISM utilizes both the prioritized online buffer and the offline dataset to train the skill encoder $q$, skill prior $p$, and low-level policy $\pi_{l}$, following the SPiRL framework (Pertsch et al., 2021). The refined low-level policy $\pi_{l}$ is shared across tasks, while the high-level policy $\pi_{h}$ leverages these refined skills to perform skill-based meta-RL within the hierarchical structure of SiMPL (Nam et al., 2022). The overall structure of PRISM is illustrated in Fig. 3. In this process, skills are iteratively refined using high-quality samples from the prioritized online buffer, enhancing the high-level policy’s ability to improve task-solving performance on both training and test tasks.
4.3 Maximum Return Relabeling for Skill Prioritization
Figure 4: An illustration of maximum return relabeling
The proposed PRISM framework trains the low-level policy $\pi_{l}$ by leveraging both the offline dataset $\mathcal{B}_{\text{off}}$ and the prioritized online buffers $\mathcal{B}_{\text{on}}^{i}$ for each training task $\mathcal{T}_{i}$. While $\mathcal{B}_{\text{on}}^{i}$ provides high-quality trajectories tailored to training tasks, relying solely on it can lead to overfitting and poor generalization. Conversely, $\mathcal{B}_{\text{off}}$ includes essential trajectories for generalization but also contains noisy and irrelevant data, as illustrated in Fig. 1. To balance these, PRISM combines both $\mathcal{B}_{\text{off}}$ and $\mathcal{B}_{\text{on}}^{i}$ to effectively train $\pi_{l}$.
However, as $\mathcal{B}_{\text{off}}$ often includes noisy and irrelevant trajectories, uniform sampling from it can unnecessarily expand the skill space and hinder the high-level policy $\pi_{h}$ from efficiently identifying task-relevant skills. To address this, PRISM employs maximum return relabelling, a method that evaluates the importance of trajectories in $\mathcal{B}_{\text{off}}$ by assigning hypothetical returns based on their task relevance. This prioritization ensures only the most useful trajectories contribute to skill learning. To compute the hypothetical return for each trajectory, PRISM trains a reward model $\hat{R}(s_{t},a_{t},i)$ for each training task $\mathcal{T}_{i}$. The reward model is optimized using the following loss function:
$\vspace{-.5em}\mathbb{E}_{(s_{t}^{i},a_{t}^{i},r_{t}^{i})\sim\mathcal{B}_{\mathrm{exp}}^{i}\cup\mathcal{B}_{\mathrm{on}}^{i}}[(\hat{R}(s^{i}_{t},a^{i}_{t},i)-r_{t}^{i})^{2}],$ (4)
where the ground truth rewards $r_{t}^{i}$ are derived from the exploration buffers $\mathcal{B}_{\text{exp}}^{i}$ and prioritized online buffers $\mathcal{B}_{\text{on}}^{i}$ for all tasks. The trained reward model is used to assign hypothetical returns $\sum_{t}\gamma^{t}\hat{R}(s_{t},a_{t},i)$, and the maximum return $\hat{G}$ for each trajectory $\tilde{\tau}\in\mathcal{B}_{\text{off}}$ is then calculated as:
$\hat{G}(\tilde{\tau}):=\max_{i}\left\{\sum_{t}\gamma^{t}\hat{R}(s_{t},a_{t},i)\right\}.\vspace{-.5em}$ (5)
Here, $\hat{G}$ captures the highest return among all training tasks for a given trajectory $\tilde{\tau}$, ensuring that only trajectories with significant potential for solving training tasks are prioritized. Using the estimated $\hat{G}$, we employ a prioritization method to select the most important trajectories for training the skill policy $\pi_{l}$, skill encoder $q$, and skill prior $p$. The loss function $\mathcal{L}_{\mathrm{skill}}(\pi_{l},q,p)$, modified from Eq. (1) as:
@@ 4.3 Maximum Return Relabeling for Skill Prioritization
$\mathcal{L}_{\mathrm{skill}}(\pi_{l},q,p):=$ $(1-\beta)\mathbb{E}_{\begin{subarray}{c}\tilde{\tau}\sim P_{\mathcal{B}_{\text{off}}},\\ z\sim q(\cdot|\tilde{\tau})\end{subarray}}\left[\mathcal{L}(\pi_{l},q,p,z)\right]$ (6)
$+\frac{\beta}{N_{\mathcal{T},\mathrm{train}}}\sum_{i}\mathbb{E}_{\begin{subarray}{c}\tau^{i}\sim\mathcal{B}_{\text{on}}^{i},\\ z\sim q(\cdot|\tau^{i})\end{subarray}}\left[\mathcal{L}(\pi_{l},q,p,z)\right],$
where $\mathcal{L}(\pi_{l},q,p,z):=-\sum^{t+H_{s}-1}_{k=t}\log\pi_{l}(a_{k}|s_{k},z)+\lambda^{\text{kld}}_{l}\mathcal{D}_{\text{KL}}\left(q||\mathcal{N}(\mathbf{0},\mathbf{I})\right)+\mathcal{D}_{\text{KL}}(\lfloor q\rfloor||p)$ is the skill loss defined in Eq. (1), and $P_{\mathcal{B}_{\text{off}}}(\tilde{\tau})=\mathrm{Softmax}(\hat{G}(\tilde{\tau})/T)$ denotes the sampling probability for prioritization, computed by applying a Softmax function over all trajectories in $\mathcal{B}_{\text{off}}$, with $T>0$ as the temperature parameter. For $\mathcal{B}_{\text{on}}$, which already contains high-return trajectories, trajectories are sampled uniformly. The control factor $\beta$ dynamically adjusts the balance between the offline and online datasets during training based on their average returns, given by
$\beta=\frac{\exp(\bar{G}_{\text{on}}/T)}{\exp(\bar{G}_{\text{on}}/T)+\exp(\bar{G}_{\text{off}}/T)},\vspace{-.5em}$ (7)
@@
where $\bar{G}_{\text{off}}$ represents the average $\hat{G}$ across all trajectories in $\mathcal{B}_{\text{off}}$, and $\bar{G}_{\text{on}}$ is the average actual return in $\mathcal{B}_{\text{on}}^{i}$ across all tasks. Fig. 4 illustrates the prioritization process, showing how $\beta$ dynamically balances contributions from offline and online datasets. This mechanism ensures the selection of task-relevant trajectories from both datasets, facilitating efficient training of the low-level policy. In Section 5, we evaluate the effectiveness of this approach in improving $\pi_{h}$ training compared to using $\mathcal{B}_{\text{off}}$ without prioritization. We also analyze the evolution of $\beta$ during training and the characteristics of trajectories selected for skill learning.
For implementation, PRISM refines skills every $K_{\text{iter}}$ iterations and reinitializes $\pi_{h}$ to avoid non-stationary issues. Before training, the low-level policy $\pi_{l}$, skill encoder $q$, and skill prior $p$ are pre-trained on the offline dataset $\mathcal{B}_{\text{off}}$ using Eq. (1). Exploration policy is trained using the soft actor-critic (SAC) algorithm (Haarnoja et al., 2018), which incorporates entropy to encourage exploration. Intrinsic rewards for exploring novel trajectories are generated via random network distillation (RND) (Burda et al., 2018), leveraging a randomly initialized target network for exploration. The meta-training process of PRISM is outlined in Algorithm 1, with additional details on the meta-test phase, meta-RL losses with value functions provided in Appendix A.
Algorithm 1 PRISM: Meta-Train Phase
0:  Training tasks $\mathcal{M}_{\text{train}}$, offline dataset $\mathcal{B}_{\text{off}}$, and low-level policy $\pi_{l}$, skill encoder $q$, and skill prior $p$.
0:  High-level policy $\pi_{h}$, exploration policy $\pi_{\text{exp}}$, task encoder $q_{e}$, and reward model $\hat{R}$.
1:  Fix $\bar{\pi}_{l}\leftarrow\pi_{l}$, $\bar{q}\leftarrow q$, and $\bar{p}\leftarrow p$.
2:  for iteration $k=1,2,~{}\cdots$ do
3:     for task $i=1$ to $N_{\mathcal{T},\text{train}}$ do
4:        Collect high-level trajectories $\tau_{h}^{i}$ and low-level trajectories $\tau_{l}^{i}$ using $\pi_{h}$ with $\bar{\pi}_{l},q_{e}$.
5:        Collect exploration trajectories $\tau_{\text{exp}}^{i}$ using $\pi_{\text{exp}}$.
6:        Filter high-return trajectories $\tau_{\text{high}}^{i}$ from $\tau_{l}^{i}$ and $\tau_{\text{exp}}^{i}$ s.t. $G>\min_{\tau^{\prime}\in\mathcal{B}_{\text{on}}^{i}}G(\tau^{\prime})$.
7:        Store $\tau_{h}^{i}$, $\tau_{\text{exp}}^{i}$, and $\tau_{\text{high}}^{i}$ into $\mathcal{B}_{h}^{i}$, $\mathcal{B}_{\text{exp}}^{i}$, and $\mathcal{B}_{\text{on}}^{i}$.
8:     end for
9:     Compute prioritization factors: $P_{\mathcal{B}_{\text{off}}}$ and $\beta$.
10:     for gradient step do
11:        Update $\pi_{h}$, $q_{e}$ using Eq. (2) (Meta-RL).
12:        Update $\pi_{\text{exp}}$ using Eq. (3) (Skill exploration).
13:        Update $\pi_{l}$, $q$, $p$ using Eq. (6) (Skill refinement).
14:        Update reward model $\hat{R}$ using Eq. (4).
15:     end for
16:     if $k\bmod K_{\text{iter}}=0$ then
17:        Update $\bar{\pi}_{l}\leftarrow\pi_{l}$, $\bar{q}\leftarrow q$, and $\bar{p}\leftarrow p$.
18:        Reinitialize $\pi_{h}$.
19:     end if
20:  end for

### v3

@@ 4.1 Motivation: Toward Robust Skill Learning under Noisy Demonstrations
Figure 2: Comparison of prior skill learning methods in microwave-opening task: (a) Learned skills with expert and noisy demonstrations. (b) Meta-RL performance with the learned skills.
Most existing skill-based meta-RL approaches discussed in Section 3 assume clean offline demonstrations, but real-world datasets are often corrupted by noise from aging hardware, disturbances, or sensor drift. Unlike online training that can adapt through continuous re-training, static offline datasets are particularly susceptible to such noise. This issue becomes critical in long-horizon tasks, where errors accumulate, and in precise manipulation tasks that require reliable execution. Fig. 2(a) illustrates this problem: in the Kitchen microwave-opening task, skills learned from expert demonstrations complete the task successfully, whereas skills learned from noisy data fail even to grasp the handle. This results in a significant downstream performance drop, as shown in Fig. 2(b), where noisy skills lead to poor task success rates and unstable training curves, with each iteration denoting a training loop consisting of policy rollout and update. The root cause is that existing methods treat all trajectories equally, allowing low-quality samples to dominate skill learning.
To address this, we propose the Self-Improving Skill Learning (SISL) framework, which enhances meta-RL by introducing a decoupled skill improvement policy. The high-level policy maximizes returns using the current skill library, while the improvement policy independently perturbs trajectories near the offline data distribution to discover higher-quality variants. The resulting trajectories are selectively stored and prioritized, then used to refine the skill encoder and low-level policy. Through this iterative refinement, SISL progressively denoises the skill library and improves generalization. We further incorporate prioritized buffering and maximum return relabeling as auxiliary mechanisms that enhance sample efficiency and accelerate convergence.
4.2 Self-Improving Skill Refinement with Decoupled Policies
We now describe the proposed Self-Improving Skill Learning (SISL) framework, which formalizes the iterative refinement process motivated in Section 4.1. We adopt the standard skill-based meta-RL setup where the skill encoder $q(z|\tilde{\tau}_{t:t+H_{s}})$ extracts a latent skill $z$ from trajectory segments, the high-level policy $\pi_{h}(z|s,e^{\mathcal{T}})$ selects $z$ every $H_{s}$ steps, and the low-level skill policy $\pi_{l}(a|s,z)$ executes the chosen skill for $H_{s}$ steps. Building on this setup, SISL decouples training into two complementary components: a high-level policy $\pi_{h}$ that exploits the current skill library to maximize return, and a skill-improvement policy $\pi_{\mathrm{imp}}(a_{t}|s_{t},i)$, defined for each training task $\mathcal{T}_{i}$ ($i=1,\dots,N_{\mathcal{T},\mathrm{train}}$), that deliberately perturbs trajectories near the offline data distribution to discover improved behaviors. This decoupling enables simultaneous exploitation and targeted skill improvement. Perturbations are guided by intrinsic motivation signals (Burda et al., 2018) to encourage coverage of diverse trajectories, although any novelty-driven exploration mechanism could be used. Importantly, unlike generic novelty-driven exploration, $\pi_{\mathrm{imp}}$ restricts perturbations to remain close to the offline data manifold, producing realistic trajectories that can be effectively used to refine $\pi_{l}$. However, when the offline dataset contains noise, training $\pi_{\mathrm{imp}}$ near this distribution in the early stages may be hindered by low-quality samples, reducing the effectiveness of skill improvement.
To overcome this issue, SISL employs a self-supervised guidance mechanism using two additional buffers: The improvement buffer $\mathcal{B}_{\mathrm{imp}}^{i}=\{\tau_{\mathrm{imp}}^{i}\}$ stores all trajectories generated by $\pi_{\mathrm{imp}}$, where each trajectory is $\tau_{\mathrm{imp}}^{i}=(s_{0}^{i},a_{0}^{i},r_{0}^{i},\dots,s_{H}^{i})$ with $a_{t}^{i}\sim\pi_{\mathrm{imp}}(\cdot|s_{t},i)$, and is directly used to update $\pi_{\mathrm{imp}}$ itself, and the prioritized online buffer $\mathcal{B}_{\mathrm{on}}^{i}=\{\tau_{\mathrm{high}}^{i}\}$ selectively retains the highest-return trajectories, where each $\tau_{\mathrm{high}}^{i}$ is chosen based on its return $G(\tau^{i})=\sum_{t=0}^{H-1}\gamma^{t}r_{t}^{i}$. This buffer is initialized with the offline dataset $\mathcal{B}_{\mathrm{off}}$ and gradually becomes dominated by improved trajectories generated by both $\pi_{\mathrm{imp}}$ and $\pi_{h}$. This buffer serves two purposes: (1) it provides a cleaner, progressively improving dataset that supervises $\pi_{\mathrm{imp}}$ in a self-supervised manner, guiding it toward regions that have empirically led to success, and (2) it supplies high-quality samples for refining the skill encoder $q$, skill prior $p$, and low-level policy $\pi_{l}$.
The skill-improvement policy is then trained so that its trajectory distribution increasingly matches the successful samples in $\mathcal{B}_{\mathrm{on}}^{i}$ while still exploring variations through controlled perturbations:
$\sum_{i}\mathbb{E}_{\tau^{i}\sim\mathcal{B}_{\mathrm{imp}}^{i}\cup\mathcal{B}_{\mathrm{on}}^{i}}\big[\mathcal{L}^{\mathrm{RL}}_{\mathrm{imp}}(\pi_{\mathrm{imp}})\big]+\lambda^{\mathrm{kld}}_{\mathrm{imp}}\mathbb{E}_{\tau^{i}\sim\mathcal{B}_{\mathrm{on}}^{i}}\mathcal{D}_{\mathrm{KL}}(\hat{\pi}_{d}^{i}\|\pi_{\mathrm{imp}}),\vskip-5.0pt$ (3)
where $\lambda^{\mathrm{kld}}_{\mathrm{imp}}$ is the KLD coefficient and $\hat{\pi}_{d}^{i}$ denotes the empirical action distribution derived from $\mathcal{B}_{\mathrm{on}}^{i}$. $\mathcal{L}^{\mathrm{RL}}_{\mathrm{imp}}$ consists of the standard RL loss combined with intrinsic reward terms for perturbation, with details provided in Appendix B. This loss encourages $\pi_{\mathrm{imp}}$ to iteratively focus on more promising state-action regions discovered so far, effectively turning $\mathcal{B}_{\mathrm{on}}^{i}$ into a self-improving curriculum that steers exploration away from noisy or uninformative samples.
Building on the improved trajectory distribution obtained from this update, SISL performs skill refinement in dedicated phases every $K_{\mathrm{iter}}$ iterations rather than after every update step. At the end of each phase, the skill encoder $q$, skill prior $p$, and low-level policy $\pi_{l}$ are re-trained on both $\mathcal{B}_{\mathrm{off}}$ and $\mathcal{B}_{\mathrm{on}}^{i}$ using the SPiRL objective in Eq. 1, and the high-level policy $\pi_{h}$ is reinitialized to fully benefit from the updated skill library. This periodic refinement mitigates bias from outdated skill embeddings and accelerates adaptation, as shown in our ablation study in Appendix G.4. These phases yield progressively denoised skills and a stable training signal, enabling $\pi_{h}$ to solve tasks more efficiently as training progresses. The overall SISL structure is illustrated in Fig. 4. However, the large number of noisy trajectories in $\mathcal{B}_{\mathrm{off}}$ can still reduce skill learning efficiency. The next section introduces a maximum return relabeling mechanism that prioritizes the most relevant offline trajectories so that only high-value samples significantly influence skill refinement.
4.3 Skill Prioritization via Maximum Return Relabeling
Figure 3: The SISL framework
Figure 4: Illustration of maximum return relabeling
The proposed SISL refines the low-level skills by leveraging both the offline dataset $\mathcal{B}_{\text{off}}$ and the prioritized online buffers $\mathcal{B}_{\text{on}}^{i}$ for each training task $\mathcal{T}_{i}$. While $\mathcal{B}_{\text{on}}^{i}$ provides high-quality trajectories collected from the skill-improvement policy $\pi_{\mathrm{imp}}$ and successful rollouts from $\pi_{h}$, relying solely on it risks overfitting to a narrow distribution and limiting generalization. Conversely, $\mathcal{B}_{\text{off}}$ provides diverse trajectories that are beneficial for generalization but also contains many noisy or suboptimal rollouts, which can degrade skill quality if sampled uniformly. To address this trade-off, we introduce skill prioritization via a maximum return relabeling mechanism that assigns hypothetical returns to offline trajectories and reweights samples in both $\mathcal{B}_{\text{off}}$ and $\mathcal{B}_{\text{on}}^{i}$ according to their estimated task relevance. Specifically, maximum return relabeling assigns each trajectory $\tilde{\tau}\in\mathcal{B}_{\text{off}}$ a hypothetical return that reflects its potential contribution to task success. To compute this return, SISL trains a reward model $\hat{R}(s_{t},a_{t},i)$ for each task $\mathcal{T}_{i}$, optimized with the regression loss
$\mathbb{E}_{(s_{t}^{i},a_{t}^{i},r_{t}^{i})\sim\mathcal{B}_{\mathrm{imp}}^{i}\cup\mathcal{B}_{\mathrm{on}}^{i}}\!\!~\big[(\hat{R}(s^{i}_{t},a^{i}_{t},i)-r_{t}^{i})^{2}\big],$ (4)
where the targets $r_{t}^{i}$ come from improved trajectories generated by $\pi_{\mathrm{imp}}$ and stored in $\mathcal{B}_{\mathrm{on}}^{i}$. Since this regression is performed online using actual environment rewards, it remains stable throughout training. Using the trained reward model, SISL computes for each $\tilde{\tau}$ the maximum return
$\hat{G}(\tilde{\tau}):=\max_{i}\left\{\sum_{t}\gamma^{t}\hat{R}(s_{t},a_{t},i)\right\},$ (5)
which represents the highest predicted cumulative reward across all training tasks. The offline trajectories are then sampled according to a softmax distribution $P_{\mathcal{B}_{\text{off}}}(\tilde{\tau})=\mathrm{Softmax}\big(\hat{G}(\tilde{\tau})/T\big),$ where $T>0$ is a temperature parameter controlling prioritization sharpness. This procedure biases sampling toward promising trajectories while suppressing noisy or irrelevant ones, resulting in a cleaner training signal for skill learning. The resulting skill learning objective becomes
@@ 4.3 Skill Prioritization via Maximum Return Relabeling
$\mathcal{L}_{\mathrm{skill}}(\pi_{l},q,p):=(1-\beta)\mathbb{E}_{\begin{subarray}{c}\tilde{\tau}\sim P_{\mathcal{B}_{\text{off}}},\\ z\sim q(\cdot|\tilde{\tau})\end{subarray}}\left[\mathcal{L}(\pi_{l},q,p,z)\right]+\frac{\beta}{N_{\mathcal{T},\mathrm{train}}}\sum_{i}\mathbb{E}_{\begin{subarray}{c}\tau^{i}\sim\mathcal{B}_{\text{on}}^{i},\\ z\sim q(\cdot|\tau^{i})\end{subarray}}\left[\mathcal{L}(\pi_{l},q,p,z)\right],$ (6)
where $\mathcal{L}(\pi_{l},q,p,z)$ is the SPiRL skill loss defined in Eq. 1. Since $\mathcal{B}_{\text{on}}$ already contains high-return trajectories, samples are drawn uniformly from this buffer. In addtion, the mixing coefficient $\beta$ is computed dynamically from the offline and online datasets based on their average returns, adaptively balancing their contributions during training as
$\beta=\frac{\exp(\bar{G}_{\text{on}}/T)}{\exp(\bar{G}_{\text{on}}/T)+\exp(\bar{G}_{\text{off}}/T)},$ (7)
@@
where $\bar{G}_{\text{off}}$ is the mean $\hat{G}$ across $\mathcal{B}_{\text{off}}$ and $\bar{G}_{\text{on}}$ is the mean return in $\mathcal{B}_{\text{on}}^{i}$. Fig. 4 illustrates the prioritization process, showing how $\beta$ dynamically balances contributions from offline and online datasets. This mechanism ensures the selection of task-relevant trajectories from both datasets, facilitating efficient training of the low-level policy. As a result, meta-training yields a refined low-level policy $\pi_{l}$ and a high-level policy $\pi_{h}$ optimized over progressively cleaner data. During meta-test, $\pi_{l}$ is frozen and $\pi_{h}$ is adapted to unseen tasks using a small number of interaction trajectories, following the other skill-based meta-RL methods. Additional implementation details for the meta-train and meta-test phases, along with the algorithm table, are provided in Appendix B.